OpenStudy (anonymous):

Evaluate 3[5+3[10+4*8]]

OpenStudy (midhun.madhu1987):

Simplify the innermost bracket first..

OpenStudy (midhun.madhu1987):

10 + 4 * 8 =

OpenStudy (anonymous):

42

OpenStudy (midhun.madhu1987):

yes.. now simplify the next bracket...

OpenStudy (anonymous):

8

OpenStudy (midhun.madhu1987):

no... 3 * 42.. right?

OpenStudy (anonymous):

126

OpenStudy (midhun.madhu1987):

yes... whats the next step?

OpenStudy (anonymous):

X 8

OpenStudy (midhun.madhu1987):

no.. +5

OpenStudy (anonymous):

131

OpenStudy (midhun.madhu1987):

yes...

OpenStudy (midhun.madhu1987):

atlast one more step...

OpenStudy (anonymous):

X3=393. Right.

OpenStudy (midhun.madhu1987):

Finally... u did it.. ;P

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Evaluate [3(5+6)+3]\7

OpenStudy (midhun.madhu1987):

make a try ... i'll help...

OpenStudy (anonymous):

(5+6=11)[3+2=5]\7

OpenStudy (midhun.madhu1987):

first do (5+6), then Multiply it with 3.

OpenStudy (midhun.madhu1987):

then?

OpenStudy (anonymous):

11*3=33+2=35/7=5

OpenStudy (midhun.madhu1987):

nope

OpenStudy (midhun.madhu1987):

33 + 3 = Then divide by 7.

OpenStudy (anonymous):

Sorry it was a 2 not a 3

OpenStudy (midhun.madhu1987):

then u were correct earlier..

OpenStudy (anonymous):

Given that f(x)=x^2+2, evaluate f(1)+f(5).